OpenStudy (anonymous):

whats the answer to p(x) = ( 2x-1)(x-3)^2 find the roots

OpenStudy (debbieg):

The roots just means the values of x which are solutions to the equation p(x)=0 So you need to solve: \(\large ( 2x-1)(x-3)^2=0\)

OpenStudy (debbieg):

Keep in mind the Zero Factor Property: If AB=0, then either A=0 or B=0.

OpenStudy (anonymous):

if \(x-3=0\) then what is \(x\)?

OpenStudy (anonymous):

x = 3?

OpenStudy (anonymous):

is that the answer?

OpenStudy (debbieg):

That is one of the roots. Do you understand how to solve the equation: \(\large ( 2x-1)(x-3)^2=0\) ??

OpenStudy (anonymous):

yes that is the answer

OpenStudy (debbieg):

That is NOT the complete answer.